# Multi-LF: A Unified Continuous Learning Framework for Real-Time DDoS Detection in Multi-Environment Networks

### v1

@@ II-C Malicious Traffic Detection in IoT Networks
IoT networks have attracted significant research interest due to their lower security mechanisms compared to traditional networks [27]. For instance, Babayigit et al. [28] propose a novel approach to IoT malicious traffic detection using multiple-domain learning. They used Edge-IIoTSet [29], WUSTL-IIoT-2021 [30], and X-IIoTID [31] datasets in the proposed approach. Furthermore, they employ an autoencoder for feature-space fusion to convert all datasets into common feature space. Their hybrid deep learning model, combining CNN and GRU, achieves up to 97.68% accuracy and improvements in transfer learning scenarios. Zhu et al. [32] present the Lightweight Knowledge Distillation Space-Time Neural Network (LKD-STNN) model to address IoT security constraints by creating a compact model using knowledge distillation and adaptive temperature function dynamics. Their model achieves over 98% accuracy on ToN-IoT [33] and IoT-23 [34] datasets.
@@
Huo et al. [35] introduce LightGuard, a lightweight malicious traffic detection model for IoT. LightGuard utilizes lightweight residual block (LRB) modules (inspired by ShuffleNetV2 [36]) and a novel ghost module for efficient feature map generation, achieving over 99.6% accuracy across diverse datasets (Edge-IIoTset [29], USTCTFC2016 [37], ToN-IoT [33] and CIC-IoT [38] datasets) while maintaining low computational complexity. Babayigit et al. [28] propose a multiple-domain learning framework to improve the reliability and generalization of DL models for Industrial IoT (IIoT) traffic classification. Their work integrates Edge-IIoTSet [29], WUSTL-IIoT-2021 [30], and X-IIoTID [31] datasets using an autoencoder for dimensionality harmonization and a modified locally linear embedding for statistical alignment. They use a hybrid DL model that combines CNNs and GRUs along with Bayesian optimization for hyperparameter tuning. Their work achieves 97.68% accuracy, 97.70% recall, 97.67% precision, and 97.68% F1-score for binary classification and improves to 97.80% accuracy and 97.79% F1-score with transfer learning.
II-D Continuous Learning for Malicious Traffic Detection
@@
Continuous learning is an approach in which a model learns continuously by incorporating new data without retraining from scratch [39]. This is especially useful for adapting to evolving data in dynamic environments. Continuous learning is also important in cybersecurity due to the evolving nature of attacks. Xu et al. [40] propose an approach that uses self-paced class incremental learning (SPCIL). SPCIL leverages network traffic data to improve class incremental learning (CIL), a deep learning technique that integrates new malware classes while preserving recognition of prior categories. SPCIL uses a loss function that combines sparse pairwise loss with sparse loss. Their experimental results demonstrate that SPCIL effectively identifies both existing and new malware classes. Compared to other incremental learning methods, SPCIL excels in performance and efficiency, with a minimal parameter count of 8.35 million, achieving accuracy rates of 89.61%, 94.74%, and 97.21% in different test scenarios. Similarly, Ajjaj et al. [41] present an approach to detect black hole attacks, an attack on the Ad hoc On-Demand Distance Vector (AODV) routing protocol. They simulate realistic VANET scenarios using the Simulation of Urban Mobility (SUMO) [42] and the Network Simulator (NS-3) [43]. They evaluate the performance of two online incremental classifiers, Adaptive Random Forest (ARF) and KNN, using metrics such as accuracy, recall, precision, and F1-score, as well as training and testing time. Results demonstrate that ARF successfully classifies and detects black hole nodes in VANETs, outperforming KNN in all performance measures.
@@
Wang et al. [44] propose an incremental learning method for small-sample data to detect malicious traffic. Their method employs a pruning strategy to identify and remove redundant network structures, dynamically reallocating these resources based on the proposed measurement method according to the difficulty of the new class. Their approach ensures that the network can learn incrementally without overconsuming storage and computing resources. Additionally, the proposed method utilizes knowledge transfer to mitigate forgetting old classes, alleviating the burden of training large parameters with limited data. Their experimental results on multiple datasets outperform established baselines in classification accuracy while using 50% less memory. In addition, Zhao et al. [45] propose Trident, a framework for detecting fine-grained unknown encrypted traffic. Trident transforms the identification of known and new classes into multiple independent one-class learning tasks. It comprises three modules: tSieve for traffic profiling, tScissors for outlier threshold determination, and tMagnifier for clustering. Their evaluations on four popular datasets show that Trident outperforms 16 other related works.
II-E Multi-Environment Malicious Traffic Detection
@@
Malicious traffic detection in M-En networks remains challenging due to the heterogeneous and complex nature of traffic patterns. M-En networks, where IoT and traditional IP-based traffic coexist, create challenges for the security system. Especially, different network types exhibit distinct traffic patterns, making it challenging to develop models that can effectively capture this diversity. To address this security concern, several studies propose different approaches to securing M-En. For example, Rustam et al. [14] introduce a system that leverages the Synthetic Data Augmentation Technique (S-DATE) and a Particle Swarm Optimizer (PSO)-based Diverse-Self Ensemble Model (D-SEM) to enhance the detection of malicious activities across diverse network environments. Their approach is validated on a composite dataset integrating InSDN [46], UNSW-NB15 [13], and IoTID-20 [12], achieving an accuracy score of 98.9%.
@@
Building on this foundation, Rustam et al. [47] present a detection approach that combines a newly developed M-En traffic dataset with S-DATE to mitigate data imbalance and improve model training efficiency. This approach enhances the detection of malicious traffic, achieving a detection rate of 99.1%. In advancing cybersecurity in M-En environments, Rustam et al. [11] develop ML models trained on AI-based traffic for the M-En dataset derived from benchmark datasets UNSW-NB15 [13] and IoTID-20 [12]. Their models counter both traditional and AI-based threats, achieving accuracy rates of 98.3% for binary classification and 96.8% for multi-class problems. Similarly, Indrasiri et al. [48] propose an approach to detect malicious traffic in M-En networks through ensemble learning. By merging UNSW-NB15 and IoTID-20 datasets and reducing the feature set using Principal Component Analysis (PCA), they develop a stacked ensemble model termed Extra Boosting Forest (EBF). EBF enhances detection performance, achieving accuracy scores of 98.5% and 98.4% for binary and multi-class classifications, respectively. In addition, Zukaib et al. [49] validate a framework designed for detecting cyberattacks in dynamic Internet of Medical Things (IoMT) networks within M-En environments. Their approach integrates Federated Learning and Meta-learning within a multi-phase architecture, achieving an accuracy of 99.82%.
TABLE I: Summary of Related Work
Ref Year ML DL Dataset Method Results Purpose Limitation
[15] 2024 - CNN, Transfer Learning CIC-DDoS2019, CSE-CIC-IDS2018, UNSW-NB15 and KDDCup’99 Adaptive architectures, CNNs 93.62%, 99.92%, 99.84%, and 98.99% DDoS Detection Works with benchmark datasets, not real-time testing.
[19] 2023 Gradient Boosting, XGBoost - CICDoS2019 Comparative study GB: 99.99%, XGBoost: 99.98%, minimal false alarms DDoS Detection Limited to CICDoS2019 dataset
[1] 2023 - DFNet Self-Collected Preference-driven, in-network enforced traffic shaping 99.93% forwarding of victim-desired traffic with minimal overhead DDoS Defense False classifications and lacks real network trace evaluations
[20] 2024 LR, SVM, NB, KNN, RF, EC, SVM-RF ANN Self-Collected SVM-RF Ensemble 99.1% DDoS Attack Detection Lack of DL models evaluation.
[22] 2024 RF, ET, DT - UNSW-NB15, CIC-IDS-2017, CIC-IDS-2018 RO, Stacking Feature Embedding, PCA Accuracy: 99.59%-99.99% depending on model and dataset Network Intrusion Detection Focuses on specific datasets, and no real-time testing.
[50] 2023 - LSTM, BiLSTM, GRU, Deep RNN CIRACIC-DoHBrw-2020-time series Two-layer network classification 99% accuracy with GRU Network Intrusion Detection Focuses on DoH traffic and did not focus on real-time diverse traffic.
[25] 2023 RF, LR, KNN - NSL-KDD CorrAUC, feature selection algorithm 99% with RF Feature Selection for Traffic Classification Works with an old dataset and no real-time testing.
[26] 2023 Adaptive Random Forests - NSL-KDD Online active learning, adaptive margin sampling 98.20% accuracy, outperforms other state-of-the-art methods Complex Traffic Detection Focuses on NSL-KDD dataset
[28] 2024 - CNN, GRU Edge-IIoTSet, WUSTL-IIoT-2021, X-IIoTID Multiple-domain learning, autoencoder, modified locally linear embedding Up to 97.68% accuracy, significant improvements in transfer learning scenarios IoT Malicious Traffic Detection Limited to specific IoT datasets
[32] 2023 - LKD-STNN ToN-IoT, IoT-23 Knowledge distillation, adaptive temperature function dynamics Over 98% accuracy IoT Security Focuses on ToN-IoT and IoT-23 datasets
[35] 2023 - - Edge-IIoTset, USTCTFC2016, ToN-IoT and CIC IoT Dataset 2023 Lightweight residual block (LRB) modules, ghost module Over 99.6% accuracy for CIC IoT Dataset 2023, 99.94% USTC-TFC2016, 99.93% ToN-IoT and 99.9% Edge-IIoTset IoT Malicious Traffic Detection Limited to specific IoT datasets and no real-time testing.
[40] 2023 - SPCIL USTC-TFC2016 Self-paced class incremental learning (SPCIL) Accuracy rates: 89.61%, 94.74%, and 97.21% in increments of 2, 4, and 5 Incremental Learning for Malicious Traffic Focuses on small-sample datasets
[41] 2023 ARF, KNN - Simulated VANET scenarios Incremental learning, Adaptive Random Forests, K-Nearest Neighbors ARF outperforms KNN in all performance measures Incremental Learning for VANET Security Higher training and testing time for ARF
[44] 2023 - - - Incremental learning, pruning strategy, knowledge transfer Superior performance, 50% less memory usage Small-sample Malicious Traffic Classification Focuses on small-sample datasets
[45] 2024 - - Four popular datasets tSieve, tScissors, tMagnifier Significantly outperforms 16 state-of-the-art methods Fine-grained Unknown Encrypted Traffic Detection Customizable framework for specific scenarios
[14] 2024 - - InSDN, UNSW-NB15, IoTID-20 Synthetic Data Augmentation Technique (S-DATE), Particle Swarm Optimizer (PSO) Accuracy score: 0.989 M-En Malicious Traffic Detection Focuses on synthetic data augmentation technique
[47] 2023 - - M-En traffic dataset S-DATE, new M-En traffic dataset Detection rate: 0.991 Mitigating Data Imbalance in M-En Networks Focuses on specific M-En traffic dataset
[11] 2024 Extra Trees - UNSW-NB15, IoTID-20 AI-based traffic detection Accuracy: 0.983 for binary, 0.968 for multi-class M-En Network Security Focuses on benchmark datasets
[48] 2022 Extra Tree, Gradient Boosting, Random Forest - UNSW-NB15, IoTID-20 Ensemble learning, PCA Accuracy: 0.985 for binary, 0.984 for multi-class Malicious Traffic Detection in M-En Networks Focuses on specific datasets
[49] 2024 - - IoMT networks Federated Learning, Meta-learning Accuracy: 99.82% Cyberattack Detection in IoMT Networks Focuses on IoMT networks
II-F The Gap in the Literature
@@
Despite significant advancements in malicious traffic detection using ML across various network environments, challenges remain. Most studies focus on specific network architectures, such as IoT [35] or traditional networks [22]. There is a lack of research in the M-En domain, with few studies [47, 49] relying on synthetic combinations of existing benchmark datasets, as no real M-En datasets exist. This reliance on synthetic datasets may limit generalizability to real-world M-En networks. Moreover, adapting to continuously evolving threats, particularly advanced persistent threats utilizing sophisticated attack strategies, presents ongoing challenges. Some studies adopt continuous learning [44], but they are domain-specific, focusing either on IoT or other networks. Therefore, an approach is needed to handle M-En networks in real-time and update their knowledge with continuous learning.
Figure 1: Overview of the proposed framework for malicious traffic detection for M-En networks
III Proposed Methodology
We propose a real-time approach for malicious traffic detection in M-En networks using ML. In this real-time approach, we utilize a continuous learning technique that enables the ML models to continuously learn and adapt to new traffic patterns, ensuring the system remains up-to-date over time. We evaluate and validate our proposed approach in realistic scenarios by generating real-world traffic and using our models to classify this traffic in real-time.
Figure 1 shows our proposed approach and its four main phases: Dataset Collection, Feature Extraction, ML Approach, and Continuous Learning.
In the first phase (Dataset Collection), we collect a PCAP dataset (a dataset that consists of recorded data packets that have traversed in a network). We collect this dataset by leveraging DDoShield-IoT [5] (a testbed that uses NS-3 and Docker containers to replicate realistic IoT environments and traffic for studying intrusion detection systems) to generate real-world malicious IoT traffic and benign traditional traffic that are supported by DDoShield-IoT. Since DDoShield-IoT does not support the generation of benign IoT traffic and malicious traditional traffic, we leverage existing PCAP datasets (published in the literature) that contain both of the traffic types that DDoShield-IoT cannot generate. In this manner, we are able to construct a dataset that contains real IoT benign and malicious traffic as well as real traditional (non-IoT) benign and malicious traffic. We discuss our Dataset Collections phase in detail in §III-A.
In the second phase (Feature Extraction), we extract general features directly from the packets in the PCAP dataset (e.g., destination ports). We also calculate statistical features for these packets (e.g., the average number of packets received in a time window). Then, we merge these general and statistical features to form a significant feature set. These statistical and general features generate a more correlated and linearly separable feature set for malicious and benign targets to train the model accurately and efficiently. We discuss the details of the feature extraction phase in §III-B.
In the third phase (ML Approach), we start by preprocessing the merged feature set. After the preprocessing step, we split the processed feature set into training and testing sets. Then, we use the training set to train multiple ML models and subsequently test these trained models on the testing set to check their performance. We estimate the performance of the ML models in terms of accuracy, precision, recall, F1 score, ROC curve, and computational time. We discuss the details of the ML Approach phase in III-C.
In the last phase (Continuous Learning), we deploy our trained ML models using DDoShield-IoT and validate their performance in real-time. In addition, we utilize a continuous learning technique to retrain the model on new (unseen) data over time. We use continuous learning in two scenarios: with and without human (expert) involvement. We discuss the details of the Continuous Learning phase in §III-D.
III-A Phase 1: Data Collection

### v2

@@ 2.3 Malicious Traffic Detection in IoT Networks
IoT networks have attracted significant research interest due to their lower security mechanisms compared to traditional networks [omolara2022internet, LEFOANE2025104110]. For instance, Babayigit et al. [babayigit2024towards] propose a novel approach to IoT malicious traffic detection using multiple-domain learning. They used Edge-IIoTSet [9751703], WUSTL-IIoT-2021 [zolanvari2021wustl], and X-IIoTID [9504604] datasets in the proposed approach. Furthermore, they employ an autoencoder for feature-space fusion to convert all datasets into a common feature space. Their hybrid deep learning model, combining CNN and GRU, achieves up to 97.68% accuracy and improvements in transfer learning scenarios. Zhu et al. [10236538] present the Lightweight Knowledge Distillation Space-Time Neural Network (LKD-STNN) model to address IoT security constraints by creating a compact model using knowledge distillation and adaptive temperature function dynamics. Their model achieves over 98% accuracy on ToN-IoT [MOUSTAFA2021102994] and IoT-23 [garcia2020iot23] datasets.
@@
Huo et al. [10536057] introduce LightGuard, a lightweight malicious traffic detection model for IoT. LightGuard utilizes lightweight residual block (LRB) modules (inspired by ShuffleNetV2 [vnwt]) and a novel ghost module for efficient feature map generation, achieving over 99.6% accuracy across diverse datasets (Edge-IIoTset [9751703], USTCTFC2016 [lu2016ustc], ToN-IoT [MOUSTAFA2021102994] and CIC-IoT [unb_iot_dataset] datasets) while maintaining low computational complexity. Babayigit et al. [babayigit2024towards] propose a multiple-domain learning framework to improve the reliability and generalization of DL models for Industrial IoT (IIoT) traffic classification. Their work integrates Edge-IIoTSet [9751703], WUSTL-IIoT-2021 [zolanvari2021wustl], and X-IIoTID [9504604] datasets using an autoencoder for dimensionality harmonization and a modified locally linear embedding for statistical alignment. They use a hybrid DL model that combines CNNs and GRUs along with Bayesian optimization for hyperparameter tuning. Their work achieves 97.68% accuracy, 97.70% recall, 97.67% precision, and 97.68% F1-score for binary classification and improves to 97.80% accuracy and 97.79% F1-score with transfer learning.
2.4 Continuous Learning for Malicious Traffic Detection
@@
Continuous learning is an approach in which a model learns continuously by incorporating new data without retraining from scratch [fan2019real]. This approach is especially useful for adapting to evolving data in dynamic environments. Continuous learning is also important in cybersecurity due to the evolving nature of attacks. Xu et al. [10472312] propose an approach that uses self-paced class incremental learning (SPCIL). SPCIL leverages network traffic data to improve class incremental learning (CIL), a deep learning technique that integrates new malware classes while preserving recognition of prior categories. SPCIL uses a loss function that combines sparse pairwise loss with sparse loss. Their experimental results demonstrate that SPCIL effectively identifies both existing and new malware classes. Compared to other incremental learning methods, SPCIL excels in performance and efficiency, with a minimal parameter count of 8.35 million, achieving accuracy rates of 89.61%, 94.74%, and 97.21% in different test scenarios. Similarly, Ajjaj et al. [ajjaj2023incremental] present an approach to detect black hole attacks, an attack on the Ad hoc On-Demand Distance Vector (AODV) routing protocol. They simulate realistic VANET scenarios using the Simulation of Urban Mobility (SUMO) [krajzewicz2010traffic] and the Network Simulator (NS-3) [riley2010ns]. They evaluate the performance of two online incremental classifiers, Adaptive Random Forest (ARF) and KNN, using metrics such as accuracy, recall, precision, and F1-score, as well as training and testing time. Results demonstrate that ARF successfully classifies and detects black hole nodes in VANETs, outperforming KNN in all performance measures.
@@
Wang et al. [wang2023drnet] propose an incremental learning method for small-sample data to detect malicious traffic. Their method employs a pruning strategy to identify and remove redundant network structures, dynamically reallocating these resources based on the proposed measurement method according to the difficulty of the new class. Their approach ensures that the network can learn incrementally without overconsuming storage and computing resources. Additionally, the proposed method utilizes knowledge transfer to mitigate forgetting old classes, alleviating the burden of training large parameters with limited data. Their experimental results on multiple datasets outperform established baselines in classification accuracy while using 50% less memory. In addition, Zhao et al. [zhao2024trident] propose Trident, a framework for detecting fine-grained unknown encrypted traffic. Trident transforms the identification of known and new classes into multiple independent one-class learning tasks. It comprises three modules: tSieve for traffic profiling, tScissors for outlier threshold determination, and tMagnifier for clustering. Their evaluations on four popular datasets show that Trident outperforms 16 other related works.
2.5 Malicious Traffic Detection in Multi-Environment Networks
@@
Malicious traffic detection in M-En networks remains challenging due to the heterogeneous and complex nature of traffic patterns. M-En networks, where IoT and traditional IP-based traffic coexist, create challenges for the security system. Especially, different network types exhibit distinct traffic patterns, making it challenging to develop models that can effectively capture this diversity. To address this security concern, several studies propose different approaches to securing M-En. For example, Rustam et al. [rustam2024malicious] introduce a system that leverages the Synthetic Data Augmentation Technique (S-DATE) and a PSO-based Diverse-Self Ensemble Model (D-SEM) to enhance the detection of malicious activities across diverse network environments. Their approach is validated on a composite dataset integrating InSDN [9187858], UNSW-NB15 [7348942], and IoTID-20 [IoTID20], achieving an accuracy of 98.9%.
@@
Building on this foundation, Rustam et al. [rustam2023securing] present a detection approach that combines a newly developed M-En traffic dataset with S-DATE to mitigate data imbalance and improve model training efficiency. This approach enhances the detection of malicious traffic, achieving a detection rate of 99.1%. In advancing cybersecurity in M-En environments, Rustam et al. [rustam2024ai] develop ML models trained on AI-based traffic for the M-En dataset derived from benchmark datasets UNSW-NB15 [7348942] and IoTID-20 [IoTID20]. Their models counter both traditional and AI-based threats, achieving accuracy rates of 98.3% for binary classification and 96.8% for multi-class problems. Similarly, Indrasiri et al. [indrasiri2022malicious] propose an approach to detect malicious traffic in M-En networks through ensemble learning. By merging UNSW-NB15 and IoTID-20 datasets and reducing the feature set using Principal Component Analysis (PCA), they develop a stacked ensemble model termed Extra Boosting Forest (EBF). EBF enhances detection performance, achieving accuracy scores of 98.5% and 98.4% for binary and multi-class classifications, respectively. In addition, Zukaib et al. [zukaib2024meta] validate a framework designed for detecting cyberattacks in dynamic Internet of Medical Things (IoMT) networks within M-En environments. Their approach integrates Federated Learning and Meta-learning within a multi-phase architecture, achieving an accuracy of 99.82%.
Table 1: Summary of Related Work
Ref Year M-En ML DL Dataset Method Results RT RE CL Early Raw
[…]
[10485168] 2023 $\otimes$ GBM, XGB - CICDoS2019 - 99.99, 99.98 $\otimes$ $\checkmark$ $\otimes$ $\otimes$ $\otimes$
[…]
[casanova2023malicious] 2023 $\otimes$ - LSTM, BiLSTM, GRU, RNN CIRACIC-DoHBrw2020 TNC 99.00 $\otimes$ $\otimes$ $\otimes$ $\otimes$ $\otimes$
[…]
[10543240] 2023 $\otimes$ ARF - NSLKDD OAL, AMS 98.20 $\otimes$ $\otimes$ $\checkmark$ $\otimes$ $\otimes$
[…]
[10472312] 2023 $\otimes$ - SPCIL USTCTFC2016 AA 89.61, 94.74, 97.21 $\otimes$ $\otimes$ $\checkmark$ $\otimes$ $\otimes$
[…]
[wang2023drnet] 2023 $\otimes$ - 1D-CNN - PS, KT 91.15 - - $\checkmark$ $\otimes$ $\checkmark$
[…]
[anley2024robust] 2024 $\otimes$ - CNN, TL CICDDoS2019, CICIDS2018, UNSWNB15, KDDCup99 AA, CNN 93.62, 99.92, 99.84, 98.99 $\otimes$ $\otimes$ $\otimes$ $\otimes$ $\otimes$
[…]
[talukder2024machine] 2024 $\otimes$ RF, ET, DT, XGB - UNSWNB15, CICIDS2017, CICIDS2018 RO, SFE, PCA 99.59, 99.99 $\otimes$ $\checkmark$ $\otimes$ $\otimes$ $\otimes$
[…]
[zhao2024trident] 2024 $\otimes$ - AE, RNN, GNN - tSieve, tScissors 96.71 $\otimes$ $\otimes$ $\checkmark$ $\otimes$ $\Box$
[…]
[rustam2024ai] 2024 $\checkmark$ ETC - UNSWNB15, IoTID20 AI-based MTC 98.30 $\otimes$ $\checkmark$ $\otimes$ $\otimes$ $\otimes$
[…]
[10948513] 2025 $\otimes$ OCSVM, IF, HBOS, INNE GAE, TE CICIDS2017, CICIDS2018, ACIIoT2023 GTAE-IDS >>98 $\Box$ $\otimes$ $\otimes$ $\checkmark$ $\checkmark$
[…]
Our 2025 $\checkmark$ SGDC, PER, MNB, BNB LSTM, LwT SC M-En AA, Multi-LF 99.99 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
  • RTE—(Real-Time Evaluation) whether the method evaluated in real-time; M-En—(Multi-Environment) whether multiple network domains are used; RE— (Resources Efficiency);whether the method is resource efficient; CL— (Continuous Learning); whether the method get update over time with recent attacks; Raw—Whether raw input data is used; Early—Whether early traffic classification is used. —– Acronyms: Transfer Learning (TL), Adaptive Architectures (AA), Self-Collected (SC), Stacking Feature Embedding (SFE), Two-layer Network Classification (TNC), Feature Selection (FS), Adaptive Random Forests (ARF), Online Active Learning (OAL), Adaptive Margin Sampling (AMS), Multiple-Domain Learning (MDL), Autoencoder (AE), Modified Locally Linear Embedding (MLLE), Knowledge Distillation (KD), Adaptive Temperature Function Dynamics (ATFD), Lightweight Residual Block (LRB), Ghost Module (GM), Self-Paced Class Incremental Learning (SPCIL), Pruning Strategy (PS), Knowledge Transfer (KT), Ensemble Learning (EL), Federated Learning (FL), Meta Learning (MeL), Light-weighted Transformer (LwT), Graph-Autoencoder (G-AE), Transformer-based Encoding (TE), Isolation-based Anomaly Detection using Nearest Neighbor Ensembles (INNE), Histogram-based Outlier Score (HBOS), Isolation forest (IF), One-class Support Vector Machines (OCSVM), —– Symbols: $\checkmark$ present, $\Box$ partial, $\otimes$ lacking; “-” indicates that the information is not provided in the paper.
2.6 Gaps & Limitations
@@
Despite significant advancements in malicious traffic detection using ML across various network environments, challenges remain. Most studies focus on specific network architectures, such as IoT [10536057] or traditional networks [talukder2024machine]. There is a lack of research in the M-En domain, with few studies [rustam2023securing, zukaib2024meta] relying on synthetic combinations of existing benchmark datasets, as no realistic M-En datasets exist. This reliance on synthetic datasets may limit the generalizability of findings to real-world M-En networks. Most studies performed M-En experiments by merging datasets from different domains [rustam2024famtds, rustam2024malicious1], which can introduce unrealistic traffic patterns and lose critical semantics due to disrupted context and feature reduction [el2009impact]. Moreover, adapting to continuously evolving threats, particularly advanced persistent threats utilizing sophisticated attack strategies, presents ongoing challenges. Some studies adopt continuous learning [wang2023drnet], but they are domain-specific, focusing either on IoT or other networks without human oversights [rustam2025adaptive]. Therefore, an approach is needed to handle M-En networks in real-time and update their knowledge with continuous learning.
3 Proposed Methodology
We introduce a real-time malicious traffic detection framework for M-En networks that combines machine learning with online, continuous learning, allowing the models to adapt to emerging traffic patterns as they emerge. Figure 1 outlines the workflow, which proceeds through four phases: (i) Dataset Collection, (ii) Feature Extraction, (iii) ML Approach, and (iv) Continuous Learning. We discuss the details of each phase in the following subsections.
Figure 1: Proposed Framework For M-En Malicious Traffic Detection
3.1 Phase 1: Data Collection